40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 22 16:30:31 2013, Koji, Update, General, Vent preparation 
    Reply  Mon Jul 22 16:56:14 2013, Jenne, Update, General, Vent preparation - In progress 
       Reply  Mon Jul 22 17:03:25 2013, Jenne, Update, General, MC spot positions MCdecenter_22July2013.png
          Reply  Tue Jul 23 04:26:54 2013, Jenne, Update, General, MC spot positions MCdecenter_22July2013_lowpower.png
       Reply  Mon Jul 22 18:49:16 2013, Jenne, Update, General, Vent preparation - In progress 
       Reply  Mon Jul 22 21:23:04 2013, manasa, Update, General, Vent preparation - In progress 
          Reply  Tue Jul 23 03:10:57 2013, Koji, Update, General, Vent preparation - In progress OSEM_RECORD.pngOPLEVS.png
          Reply  Tue Jul 23 03:43:32 2013, manasa, Update, General, Vent preparation - In progress 
             Reply  Tue Jul 23 04:22:03 2013, Koji, Update, General, Vent preparation - In progress 
       Reply  Tue Jul 23 04:31:36 2013, Jenne, Update, General, Vent preparation - Ready for Vent 
          Reply  Tue Jul 23 08:21:31 2013, Steve, Update, VAC, Vent has started beforVent.pngstartVent.png
             Reply  Tue Jul 23 14:00:03 2013, Steve, Update, VAC, the vent is completed Atm.png5cyl.pngventSum1d.png
Message ID: 8889     Entry time: Mon Jul 22 16:30:31 2013     Reply to this: 8890
Author: Koji 
Type: Update 
Category: General 
Subject: Vent preparation 

[Annalisa, Manasa, Jenne, Koji]

We are working on the vent preparation.

First of all, there was no light in the interferometer.
Obviously there were lots of IFO activity in the weekend. Some were elogged, some were not.
Annalisa took her responsibility to restore the alignment and the arms recovered their flashes.

The odd thing was that the ASS got instable after we turned down the TRY PD gain from +20dB to +10dB (0dB original).
We increased the TRY gain by factor of 10 (that's the "10dB" of this PDA520. See the spec sheet) to compensate this change.
This made the ASS instable. Anyway we reduced the gain of TRY PD to 0dB. This restored the ASS.

Jenne took some more data for the QPD spectrum calibration.

Link to the vent plan

ELOG V3.1.3-